OpenStudy (bbb911):

LaDonna is running for city council. She needs to get 5/6 of her votes from senior citizens and 36,000 votes in all to win. How many votes does she need from voters other than the senior citizens? how many other voters does she need?

OpenStudy (lyrae):

Okay, so LaDonna need 5/6 of her votes from seniors: 5/6 * 36,000 = 30,000 The other votes needed is the differnce between the total amount and those, so: 36,000 - 30,000 = 6000 You could also do: (1 - 5/6) * 36,000 = 6000

OpenStudy (bbb911):

thank you :)

OpenStudy (lyrae):

Np ^^